OpenStudy (calculusxy):

7 + n 1 ------ = ---- 43 + n 3 Find the value of n @mathstudent55

OpenStudy (cwrw238):

you can cross multiply here 3(7 + n) = 43 + n now its straight forward to solve for n

OpenStudy (calculusxy):

Is it like solving for n?

OpenStudy (cwrw238):

yes

OpenStudy (calculusxy):

So would that be like using the Distributive Property because you distribute 3 to 7 and n which would equal to 21+n

OpenStudy (anonymous):

3(7+n)=43+n (==) 21+3n=43+; (==)3n-n=43-21(==)2n=22 donc n=11